Instrument Log 

25 M IRT C1 60 MT, C1 Lamont, OK

Date

Technician

Comments

6/20/16

JEM

Maintenance was being performed on the 25M T/H from 1715 GMT to 1730 GMT. Data may be effected during this time. RTS @1730 GMT.

6/14/16 6/15/16

CAW/TMG

SE tower booms needed to be lowered to install new cameras on tower. Lowered SE booms at 1800 GMT on 6/14. Booms remained down overnight while working on camera mounts. Raised booms after cameras were installed and RTS at 2025 GMT on 6/15.

12/3/15

JEM

Instrument will be effected while maintenance is being performed on the 25M MFR. Effected time period is 1520 GMT to 1605 GMT.

10/27/15

JEM

The SE boom was lowered @1522 GMT for swapping the UIR and US. IRT data will be effected during this time. Also cleaned the IRT while it was accessible. Raised the boom and RTS @1620 GMT.

10/26/15

CAM

Lowered the SE booms at 1435 GMT for T/H calibration changes. Raised booms and RTS at 1949 GMT. No problem with this instrument.

8/25/15

JEM

SE tower booms were lowered from 1835 GMT to 1850 GMT to work on the 25M MFR. IRT data will be effecting during this time period only.

4/9/2015

PD

PGS shed at the base of the 60M tower is being replaced with a 20' ISO container. Crane and other equipment used to remove the PGS shed may have been in the field of vision of this instrument between 1420 and 1610 GMT. All equipment was out of the field of vision by 1610 GMT.

2/2/15

JEM

The 25M MFR needs to be reinstalled. 25M IRT data will be affected from 2048 GMT to 2100 GMT. RTS @2100 GMT.

1/30/15

JEM

The boom was lowered for 25M MFR maintenance. Data will be affected from 1932 GMT to 1952 GMT. RTS @1952 GMT.

1/27/15

JEM

The boom was lowered for MFR25M maintenance. Data will be affected from 2105 GMT to 2120 GMT.

10/28/14

CAM

Lowered the SE booms at 1520 GMT for T/H calibration changes. Raised booms and RTS at 2230 GMT. No problem with this instrument.

8/27/14

JEM

The 25M US needs to be replaced for annual calibration. The East boom was lowered from 1330 GMT to 1348 GMT to replace the 25M US. Data will be affected during this period. RTS @1350 GMT.

6/23/14

CAM

No data had been collected from 25IRT since 1203 GMT on 6/21. Found the UPS that powers the comms equipment in the AOS trailer was down. Power cycled UPS and restored power at 1220 GMT. Normal data collections resumed at next interval.

5/27/14

CAM

No data had been collected from instrument since 1203 GMT on 5/25. Found the UPS that powers the comms equipment in the AOS trailer was down. Breaker for UPS supply line was tripped. Reset breaker at 1325 GMT. UPS powered up and comms equipment power restored. Normal data collections resumed at next interval.

9/19/13

JEM

The 25M IRT needs to be replaced with a newer style sensor. Also DQPR 3834 indicates a bias between the 10M and 25M IRT. R/R the *25M IRT @1430 GMT. Updated the datalogger program and RTS @1500 GMT.

*Old - 1026 (Model - KT 19.85) *New - 3354 (Model - KT 19.85 II)

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Power was restored at 1920 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/8/13

JEM

DQPR #3616 requests the 25M IRT lens be replaced. Lowered the SE booms @1942 GMT. R/R the IRT lens @1948 GMT. Raised the SE booms and RTS @1957 GMT.

12/24/12

JEM

Tower booms need to be lowered to clean the 25M IRT lens and verify IRT settings. Tower was lowered @1544 GMT. Cleaned the IRT lens. Settings were verified as expected. The insulation/housing is about to fall off of the IRT because the screws fell out. Reattached IRT insulation/housing using larger screws. Raised the tower @1627 GMT. Status OK. RTS @1630 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County. Power was restored at 2115 GMT after electrical upgrades were complete.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound. Power was restored at approximately 1930 GMT after all transformers were replaced. Tower instruments resumed normal operation.

11/5/12

CAW

Lowered booms at 1338 GMT for 25 MFR 90-day calibration. Raised booms and RTS at 1349 GMT after MFR head removed. Lowered booms again for reinstallation at 2025 GMT. Raised booms and RTS at 2036 GMT.

10/31/12

JEM

The boom needs to be lowered to verify the 25M IRT serial number. Booms were lowered at 1513 GMT. Serial number was verified as 1026. Datalogger program has 1029. Edited the program to have the proper serial number (1026) and loaded the program @1527 GMT. The booms were raised @1525 GMT.

10/9/2012

JEM

DQPR #3462 requests that t.low and t.end be adjusted on the IRT as well as UWIRT_OS be adjusted in the program to accommodate higher surface temperatures in the summer time. The IRT was lowered for unrelated maintenance on the T/H sensors. While the booms were lowered, changed t.low to 233.2 and t.end to 333.2. Also edited the program and changed UWIRT_OS to 233.2. Loaded the new program @2100 GMT. Verified readings. The booms were raised @2145 GMT. RTS @2145 GMT.

10/9/12

CAM

Lowered the SE booms at 1900 GMT for T/H calibration changes. Raised booms and RTS at 2145 GMT. No problem with this instrument.

7/13/2012

JEM

The SE tower elevator was lowered to perform work on the 25M US. Lowered the SE tower elevator @1943 GMT and raised it @ 2002 GMT.

6/26/12

CAM/PAD

Last SDS collection was on 6/25/12 at 1200 GMT. Collection log shows no response from CDL. Unable to connect to datalogger remotely. Powercycled 25M side A of tower at 1726 GMT. No change in status. Checked comms to base of tower with no problems found. Lowered side A booms at 1837 GMT. Troubleshooting found Campbell Scientific SC32a interface located at datalogger had failed. R/R interface and was able to connect to datalogger from base of tower. Raised side A booms. RTS at 2007 GMT.

6/4/12

JEM

25M IRT lowered so 60M T/RH troubleshooting can be performed. Lowered the SE booms at 1410 GMT. Raised the SE booms @1637 GMT and RTS @1640 GMT.

6/1/12

JEM

25M IRT lowered so 60M T/RH sensor can be replaced. Lowered the SE booms at 1412 GMT. Raised the SE booms and RTS @1512 GMT.

5/24/12

PAD

DQPR 3407 states there is a low bias for 60 T/H probe on side A of tower. Mentor suspects bad connection and suggests lowering/raising tower. Lowered side A 25 and 60m booms at 2115 GMT. All indicator lights for signal and AC were lit after raising booms. No issues known with instrument at this time. RTS at 2130 GMT.

5/8/12

CAW

Lowered booms at 1258 GMT for 25 MFR 90-day calibration. Raised booms and RTS at 1310 GMT after MFR head removed. Lowered booms again for reinstallation at 1522 GMT. Raised booms and RTS at 1535 GMT.

3/8/12

PAD

Radon Gas Analyzer inlet pipe has separated approximately 40M above the ground. Contractor is on site today to climb tower and repair inlet tube. Power to tower needs to be shut down before contractor can climb. Shut down power at base of tower to 25 and 60M levels, sides A and B at 2023 GMT. After repairs were made and complete inspection of inlet pipe was performed, power was restored. No issues with this instrument are known at this time. RTS at 2257 GMT.

2/8/12

JEM

The 25M MFR is scheduled for an upgrade. Tower was lowered from 1855 GMT to 2212 GMT. 25M IRT will be affected during this time.

11/1/11

CAM

Lowered the SE booms at 1815 GMT for T/H calibration changes. Raised booms and RTS at 2130 GMT. No problem with this instrument.

7/19/11

CAW

25M MFR is due for calibration. Side A tower booms were lowered at 1910 GMT to remove 25M MFR. Tower was raised at 2107 GMT. No issues known with this instrument at this time. RTS at 2110 GMT.

4/15/11

JEM

The instrument has not collected since the 0400 hour. Due to other instruments that are down the source of the problems seems to be the Rocketport hub. The Rocketport power light is flashing. Power cycling the Rocketport hub did not correct the issue. Notified the IT department of the issue.

4/7/11

CAW

MFR is due for calibration. Lowered tower at 1200 GMT and raised tower at 1215 GMT. Calibrated instrument. Lowered tower at 1433 GMT, installed instrument, and raised tower at 1444 GMT. Instrument only affected by lowering tower to calibrate 25M MFR.

11/23/10

CAW

Lowered the SE booms at 1825 GMT for 25MFR calibration. Raised booms and RTS at 2010 GMT. No problem with this instrument.

10/19/10

CAM

Tower Systems on site for tower inspections. Need to secure power for maintenance. Secured AC power to 25m and 60m levels at 1407 GMT. Restored power following inspection at 1620 GMT.

10/18/10

CAM

Lowered the SE booms at 1445 GMT for T/H calibration changes. Raised booms and RTS at 2013 GMT. No problem with this instrument.

8/19/2010

JEM

The 25IRT was disrupted during maintenance on the 25MFR from 1430 GMT to 1500 GMT.

8/18/2010

JEM

The 25IRT was disrupted during maintenance on the 25MFR from 1826 GMT to 1841 GMT.

7/14/10

CAM/JEM

60M tower SE booms need to be lowered to remove GPS array repeater. Lowered SE booms at 1445 GMT. Removed GPS equipment. Raised booms and RTS at 1556 GMT.

5/11/10

CAW

No problem with instrument at this time. 25M MFR is due for calibration. Tower was lowered at 1848 GMT to remove 25M MFR and raised at 1900 GMT. Tower was lowered again at 2040 GMT to install 25M MFR and raised at 2052 GMT. RTS at 2055 GMT.

2/16/10

CAM

25M MFR needs to be removed for calibration. There is no problem with the IRT. Lowered SE booms at 1455 GMT and removed MFR. Raised booms at 1507 GMT. Lowered booms again at 1712 GMT to reinstall MFR after it was calibrated. Raised booms and RTS at 1722 GMT.

11/19/09

CAM

SDS instrument status page indicated no data had been collected since 2103 GMT on 11/18. T/S found the UPS in the AOS trailer which powers the tower communications equipment had no output. AC LINE and ALARM LEDs were lit on the UPS. Power cycled UPS at 1342 GMT and power was restored. RTS at 1345 GMT.

11/9/09

CAM

25M MFR needs to be removed for calibration. Lowered SE booms at 1405 GMT and removed MFR. Raised booms at 1414 GMT. Lowered booms again at 1623 GMT to reinstall MFR after it was calibrated. Raised booms and RTS at 1633 GMT.

9/21/09

CAM

Lowered the SE booms at 1450 GMT for T/H calibration changes. Raised booms and RTS at 1735 GMT. No problem with this instrument.

8/25/09

CAW

SE booms lowered from 1320 to 1330 GMT to remove MFR sensor for maintenance.

5/19/09

CAM

SE booms lowered from 1743 GMT to 1758 GMT for 25m MFR maintenance. No problem with this instrument.

5/13/09

CAM

25M MFR needs to be reinstalled after it was calibrated. Lowered booms at 2146 GMT and installed MFR. Raised booms and RTS at 2201 GMT.

5/13/09

CAM

25M MFR needs to be removed for calibration. Lowered booms at 1431 GMT and removed MFR. Raised booms and RTS at 1446 GMT.

10/1/08

CAM

DQPR 2032 indicates the IRT is exhibiting a positive bias compared to the PIR. Mentor recommends cleaning lens. Lowered SE booms at 1850 GMT. Cleaned significant amount of dust build up from IRT lens. Raised booms and RTS at 1905 GMT.

7/26/07

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1333 GMT for maintenance. Raised booms and RTS at 1920 GMT.

10/11/06

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations. Lowered booms at 1447 GMT. Raised booms and RTS at 2141 GMT.

5/31/06

PAD

Jane McKinney request 60m tower lowered to record model and serial numbers from the wireless data transceiver located on SE 60M boom. Lowered side A of tower at 1755 GMT. Raised tower at 1851 GMT. No problem with this instrument. RTS at 1855 GMT.

3/31/06

PAD

Last SDS collection was at 0803 GMT. At 1453 GMT, connected laptop with instrument from AOS trailer. Instrument is active and data can be monitored. R/R Comtrol *Rocket port located in AOS trailer. RTS at 1513 GMT. *Old - 8260-002409 *New - 8260-002737

3/28/06

PAD

Could not access DSVIEW page. Don't know what time SDS first failed to collect data. Connected laptop with instrument with communication lines in AOS trailer. Could communicate and monitor instrument from there. Power cycled rocket port hub located in AOS trailer. SDS can now communicate with instrument. RTS at 1653 GMT.

3/27/06

CAM

SDS instrument status page indicated no files had been received from the 25 IRT since 1403 GMT on 3/25. Message reported is 'The CDL Has Stopped Responding'. Disconnected the fibers and bit driver from the Rocket Port hub in the AOS trailer and successfully monitored the IRT from that point with a laptop. Reconnected and had the computer department force a collection, which still failed. Power cycled the Rocket Port hub at 1925 GMT. Computer department then forced a successful collection at 1930 GMT. The following scheduled collection was successful at 2003 GMT.

2/14/06

MK

The CR10X datalogger needs to have the OS upgraded and the new upgraded CDL program installed. Lowered tower at 2000 GMT, loaded new OS on CR10X and loaded new DLD program. All readings returned to normal. R/R US and UIR desiccant and cleaned domes. Raised tower at 2050 GMT. CR10X S/N is 4501.

1/19/06

MK

The 25M IRT has stopped communicating. Found a faulty optical bit driver on the Rocket port located in the AOS trailer. Rebooted Rocket port at 1615 GMT. All communications were restored.

12/12/05

MK

The OS on all CR10X's needs to be updated. Downloaded current data off of data logger, loaded new OS on the 25M IRT system at 1945 GMT. S/N 4503. RTS at 2000 GMT.

8/5/05

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations. Lowered booms at 1327 GMT. Raised booms and RTS at 1637 GMT.

8/4/05

MR

The 25M IRT lens needs to be cleaned. Cleaned the 25M IRT lens. RTS at 1915 GMT.

8/4/05

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations. Tower Systems techs also on site for tower inspection. Turned the 25 and 60M levels power off at 1330 GMT so Tower Systems could service tower. Lowered booms at 1830 GMT. Raised booms and RTS at 2014 GMT.

6/10/05

MR

The 25m IRT has stopped sending data to the SDS. T/S indicated that the IRT was not running. Tried to reload the program but this did not fix the problem. Lowered the tower at 1945 GMT. Power cycled the data logger but the data logger would not restart. R/R the *data logger at 1950 GMT. Reloaded the program at 1953 GMT. Data logger started collecting data at 1955 GMT. Raised the tower at 2045 GMT. RTS at 2100 GMT. *Old - X26909 *New - X4501

9/30/04

JEM

Collected data from the datalogger @2010 GMT due to a collection problem with the SDS. RTS @ GMT.

9/27/04

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations. Lowered booms at 1438 GMT. Raised booms and RTS at 2056 GMT.

8/5/04

MR

The SE elevator needs to be lowered due to work on other instruments. Lowered the tower at 1410 GMT. Raised tower at 1445 GMT. RTS at 1450 GMT.

7/21/04

MR

The tower South east needs to be lowered and the instruments need to be cleaned due to the farmer working the ground under the tower. Lowered the elevator at 1526 GMT. Cleaned the IRT lens at 1600 GMT. Raised the elevator at 1615 GMT. RTS at 1626 GMT.

7/6/04

PAD

Need to power cycle SE instruments at 25M level due to communication problems with 25M MFR. There is no failure with this instrument. Power cycled 25M instruments on side A (S.E. side) of tower at 1419 GMT. RTS at 1434 GMT.

4/19/04

MR

The 25M IRT is not able to communicate with the SDS system. T/S found a bad bit driver in the comms box at the base of the tower. R/R the bit driver. RTS at 1825 GMT.

4/16/04

CAM

SDS comlogs indicated no files had been received from the 25IRT since 1502 GMT. Power cycled the SE 25M level power at 2035 GMT. Power cycled 25IRT bit driver at 2040 GMT. SDS personnel still unable to receive reply from datalogger manually. T/S continues.

4/15/04

CM/PD

SE booms need to be lowered for 25M MFR maintenance. Lowered booms at 2104 GMT. Raised booms and RTS at 2151 GMT.

2/23/04

MR

The SE tower connections need to be reseated for other equipment. Lowered the elevator about 1 foot and the raised the elevator. RTS at 1853 GMT.

2/20/04

MR

The SE elevator needs to be lowered due to work on other equipment. Lowered the elevator at 1445 GMT. Raised the elevator at 1515 GMT. RTS at 1515 GMT.

2/13/04

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the elevator at 1752 GMT. Raised the elevator at 2134 GMT. RTS at 2140 GMT.

2/10/04

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the elevator at 1905 GMT. Raised the elevator at 2315 GMT. RTS at 2330 GMT.

1/15/04

MR

The program needs to be changed and uploaded so that it reflects the new UIR calibration coefficient changes. Uploaded the new program into the data logger at 2052 GMT. RTS at 2105 GMT.

12/8/03

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1940. Raised the tower at 2105. RTS at 2105.

11/19/03

MR

The 25M and 60 M tower elevator needs to be lowered due to the IRT lens needs to be cleaned. Lowered the SE tower at 2105. Cleaned the IRT lens. Raised the tower at 2120 GMT.

10/7/03

MR

The tower connections need to be reset. Lowered the tower about 3 inches at 1530. Raised the tower at 1531. RTS at 1545.

10/6/03

MR

The tower needs to be lowered due to maintenance on other instruments. Lowered the tower at 1713 GMT. Raised the tower at 1826. RTS at 1828 GMT.

9/26/03

JEM

The 25M IRT has not collected since the 1500 GMT collection. Checked fiber optics and bit drivers. One bit driver was not functioning. Power cycled the bit driver and it began to operate normally. RTS @2100 GMT.

8/19/03

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the tower at 1415 GMT. Raised the tower at 1603. RTS at 1615.

8/15/03

CAM

Need to lower SE booms for annual 25 and 60M T/H change outs. Lowered the SE booms at 1341 GMT for maintenance. Raised booms and RTS at 1920 GMT.

8/14/03

CM

SE booms lowered a few inches at 1900 GMT to tension the cable between the ground and 25M boom. Raised booms at 1914 GMT. Only the 60M boom made connection at that time due to the added tension. Lowered booms again at 1925 GMT and repositioned the 25M boom higher on the cable. Raised booms and RTS both 25 and 60M levels at 1935 GMT.

8/14/03

CM

SE boom lowered for tower system maintenance. There was no failure with this instrument. Lowered boom at 1443 GMT. Raised boom and RTS at 1515 GMT.

8/5/03

MR

The 25M IRT has not collected since 8/1/03. T/S found a bad connection in the 25 pin data connector that connects to the bit driver. Resoldered the connection and put a new back shield on the connector due to the old one being cracked and pieces missing. RTS at 1930 GMT.

8/1/03

MR

The IRT has had intermittent collection problems with the SDS. T/S found a possible bad bit driver at the base of the tower. R/R the bit driver. RTS at 2030 GMT.

7/30/03

MR

The 25MIRT is sending bad characters when the SDS tries to collect data from it indicating a bad bit driver. R/R the bit driver in the AOS trailer. RTS at 1545 GMT.

5/21/03

JEM

Lines 44 and 47 in the 25M IRT program need to be edited to show 20 sec. and 40 sec. respectively. Edlogged program IRT25A1 and changed line 44 from 21 sec. to 20 sec. and changed line 47 from 41 sec. to 40 sec. Saved the new program as IRT25A2. Uploaded the new program to the 25M IRT @1807 GMT. Monitored the data locations and all values were reading as expected. RTS @1815 GMT.

5/19/03

JEM

The mentor has requested that the program be downloaded from the 25M IRT. Connected to the 25M IRT @1905 GMT. Downloaded the program and emailed it to the mentor. RTS @1930 GMT.

5/16/03

JEM

Instruments on 25M East boom were power cycled @2047 GMT due to problem with 25M MFR. No problems with this instrument. RTS @2100 GMT.

3/21/03

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the tower at 1824 GMT. Raised the tower at 2024 GMT. RTS at 2024 GMT.

3/18/03

JEM

The 25M IRT has not collected since 3/17/03 @1502 GMT. The power strip that powers the bit driver in the AOS trailer was found powered off. Powered the power strip back on @1505 GMT. The power strip was likely powered off by accident while rearranging the AOS system. RTS @1515 GMT.

3/17/03

MR

The instrument mentor wants the IRT and US/UIR connections checked. Checked the connections. They checked good. Due to approaching thunderstorm we had to raise the tower. Will continue with this CM when weather permits. Raised the tower at 1705 GMT.

2/21/03

MR

The 25M IRT reading has been reading about 10 degrees higher then normal. The mentor has requested that the program be reloaded. Lowered the tower at 1953 GMT. Reloaded the program at 2040 GMT. Raised the tower at 2056 GMT. RTS at 2056 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT. Power was restored to the Central Facility at approximately 1945 GMT. RTS @2000 GMT.

1/30/03

JEM

The 60M Southeast temperature has been too high. There is no problem with the 25M IRT. Corrective maintenance is being done on the 60M T/H. Lowered Southeast 25M and 60M booms @1655 GMT. RTS @1810 GMT.

12/16/02

MR

The serial numbers and calibration factors in the program for the UIR and US need to be checked and changed if necessary. Lowered the tower at 1639 GMT. The program was retrieved, checked and the calibration factors where changed. Raised the tower at 1705 GMT. RTS at 1710.

11/7/02

CM/MR

Continued T/S of communications problem. Lowered the SE booms at 1557 GMT. Loopback tests were good all the way from the Rocket Port hub in the AOS trailer to the SC32A just before the datalogger. The SC32A was also tested and found to be good, but still could not monitor the datalogger through the same communications link. Further T/S found that the MOV across the transmit and receive copper lines in the tower DC signal box was bad. With the MOV in place, communications with the IRT datalogger could not be established. But when the MOV was removed the communications were good. Tested the MOV with a DVM and it measured good as an open circuit. Replaced the MOV with a spare from a pair of lines that was not being used and communications were lost again. The only way to communicate with the datalogger is with the MOV removed from the circuit. Removed the MOV and raised the SE booms at 1755 GMT. Verified communications and RTS status O.K. at 1800 GMT. Will discuss the MOV situation with instrument mentor.

11/6/02

CM/MR

SDS comlog reports indicated that no files had been received from the 25M IRT since 1002 GMT on 11/4. T/S found that the datalogger could not be monitored through the Rocket Port hub in the AOS trailer. Lowered the SE booms at 1917 GMT for further T/S. The datalogger could be monitored by connecting a laptop directly to the datalogger through an SC32A. Checked the fiber optics at the tower base and found rodents had chewed through one cable. Replaced the fibers and the communications then tested good from the hub to the tower base, but still could not monitor the datalogger through the entire communications line. T/S isolated the problem to somewhere between the DC signal box at the base tower and the datalogger. Due to time constraints, the booms were raised at 2223 GMT and T/S will continue tomorrow.

10/31/02

CM/MR

SE booms lowered from 1445 to 1618 GMT for 25MFR and 60 T/H maintenance.

10/29/02

CAM

SE booms lowered from 2135 GMT to 2220 GMT for 25 and 60M T/H maintenance.

10/28/02

CAM

The SE 25 and 60M T/H sensors are reading 9999s or 0s. Lowered the SE booms slightly at 1731 GMT and then raised to reseat connections. RTS at 1733 GMT.

10/25/02

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1805 GMT. Raised the tower at 1912. RTS at 1920 GMT.

10/24/02

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1415 GMT. Raised the tower at 1430 GMT.

9/23/02

MR

The elevator needs to be lowered due to work being performed on the 25M MFRSR. Lowered the tower at 1930 GMT. Raised the tower at 1943 GMT. RTS at 1945 GMT.

9/10/02

MR

The instrument mentor wants the tower lowered so work can be done on the PGS sampling system. Lowered the tower at 1733 GMT. Raised tower at 1837 GMT. Status ok. RTS at 1848.

9/3/02

MR

The 25M elevator needs to be lowered to check an instrument for the PGS system. Lowered the tower at 1940 GMT. Raised the tower at 2030. RTS at 2040.

8/2/02

D. Cook

Booms lowered from 1354 to 1912 GMT for tower maintenance and 25/60m T/H calibrations.

8/1/02

D. Cook

TSI on site. Booms lowered from 1300 to 2130 GMT for tower maintenance.

5/14/02

MR

The instrument data needs to be collected and placed on the door step for the SDS to collect. Collected the data from 1519-1539 GMT. RTS at 1549 GMT.

4/11/02

CAM/MR

Lowered the SE booms at 2003 GMT for 25M MFR maintenance. Raised the booms and RTS at 2106 GMT.

4/2/02

CAM

The SE booms needed to be lowered momentarily to reseat the MFR connections. Lowered the SE booms at 1754 GMT. Raised the booms and RTS at 1755 GMT.

2/13/02

MR

The tower needs to be lowered for maintenance on other equipment. Lowered the tower at 1500. Raised the tower at 1505. RTS at 1515 GMT.

12/19/01

MR

The tower needs to be lowered due to calibration on the PGS system. Lowered the tower at 2055 GMT. Raised the tower at 2110 GMT.

11/01/01

MR

Vic Morris requested that the upwelling wiring be checked due to the readings on the US and UIR are bad. Lowered the tower at 2050 GMT. T/S found that the 100K precision resistor between the CR10 and multiplexer was not installed. Installed the resistor and notified the mentor of the change. Raised the tower at 2103 GMT.

10/11/01

MR

The SDS is ready to collect from the 25M IRT. While trying to collect from the instrument the SDS reported that the communications were lost. T/S confirmed that the communications would consist of a random C and what looked like a time stamp every few seconds. Further T/S found that the bit driver at the base of the 60M tower was bad. R/R the bit driver. This instrument needs the bit driver at the base of the tower set to DCE and the corresponding bit driver on the other end set to DTE so that communications will work. RTS at 2045 GMT.

10/4/01

MR

Joe Michalsky wants to run the calibrations on the 25M MFR a second time. Lowered the east side of the tower at 1828 GMT. Raised the tower at 1900. RTS at 1905 GMT.

10/3/01

MR

Lowered the east side of the tower at 2006 GMT for 25M MFR calibration. Raised the east side of the tower at 2057. RTS at 2105.

10/2/01

MR

The tower had to be lowered due to other maintenance. Lowered the tower at 1548. Raised the tower at 1600. RTS at 1603.

9/27/01

MR

The IRT, UIR, US is being separated from the MFRSR. Lowered the 25 and 60 meter tower at 1503. Removed the UIR, IRT and US from the MFRSR board and installed them on their own CSI data logger at 1600 GMT. Reconfigured IRT and started CSI data logger at 1627 GMT. Cleaned IRT lens at 1630 GMT. Raised the tower at 1912.

8/15/01

CAW

SE booms lowered from 1530 to 1555 GMT to R/R 25M US radiometer.

8/7/01-8/8/01

MR

Southeast booms lowered at 1918 GMT on 8/7 for 25M and 60M T/H calibrations. Booms were left down at ground level overnight. Raised booms and RTS at 1840 GMT on 8/8.

8/7/01

MR

Tower systems is on site to perform maintenance on the 25M elevator and the 60M tower. Lowered the south east 25M elevator. Tower systems replaced one of the connectors on the tower and performed general maintenance on the elevator and tower components. RTS at 1910 GMT.

7/20/01

MR

The tower will be lowered due to work being performed on the PGS system. No work will be performed on this instrument. Lowered tower at 1450 GMT. Performed PGS system maintenance. Raised tower at 1700 GMT.

7/13/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2000 GMT. Marc Fisher performed maintenance on the PGS system. Raised the tower at 2130. RTS at 2145 GMT.

7/13/01

MR

The SDS is unable to talk to the 25M T/H and 60M T/H southeast. Lowered the tower to power cycle the instruments at 1440 GMT. Raised the tower at 1441 GMT. Communications were restored.

7/12/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2100 GMT. Marc Fisher performed maintenance on the PGS system. Raised the tower at 2135. RTS at 2145 GMT.

5/18/01

MR

The instrument mentor requests that the data logger be checked for the mux chips due to the instrument reporting 00's for the IRT and upwelling radiometers. Lowered the tower at 1707 GMT. Checked for the MUX chips and observed they where missing. Installed the chips. Raised the tower and RTS at 1719 GMT.

5/4/01

MR

The data logger needs to be replaced for the system upgrade. Lowered the tower elevator at 1839 GMT. R/R the *data logger. Raised the tower and RTS at 1854 GMT. *Old - 93957 *New - 93959

11/27/00

CAW

Lowered the SE 25m and 60m booms from 1840 to 2035 GMT to clean the chilled mirrors.

9/21/2000

CAM

David Cook requested that the SE 25 and 60 meter booms be lowered and reseated as part of troubleshooting bad humidity sensor readings. Lowered the booms slightly at 1850 GMT. Raised the booms and RTS at 1852 GMT.

1/6/2000

CM/CW

25M and 60M booms lowered from 1941 GMT to 2123 GMT for installation of new GPS antenna.

11/16/99

CM/SS

The 25M IRT was insulated on 11/15/99 at mentors request. Due to time constraints, the aluminum shield could not be reformed and installed at that time. The shield will provide the sensor and insulation further protection from the elements. Lowered the 60M tower A side booms at 1700 GMT on 11/16. Removed the IRT and its mounting assembly to the lab. Reformed the aluminum shield and fitted it to the IRT mounting assembly. Replaced the IRT and raised the tower booms at 1820 GMT. RTS status O.K.

11/15/99

CM/GS

Instrument mentor requested that Site Ops insulate the 25M IRT similar to the way the 10M IRT and MWR IRT were done last year. This will improve the IRT's performance in cold weather. Lowered the 25M IRT boom at 2011 GMT. Removed the IRT to the lab to fabricate an insulation cover. Used 5/8" insulation material and aluminum tape to cover the 5 exposed sides of the IRT. Remounted the IRT, raised the tower booms and RTS at 2217 GMT. Will need to lower the boom again tomorrow to reform and replace the original aluminum shield over the insulation.

12/8/98

CM/GS

Victor Morris requested that Site Ops insulate the IRT housing to prevent data corruption during freezing temperatures. Lowered the 60m tower side A booms at 2105 GMT to take measurements. Raised the booms and RTS at 2153 GMT. Note that only measurement were made at this time. The insulation has not yet been installed.

8/6/98

GRS

Lowered the 60m tower booms on side A at 1920 GMT to read some program values on the 25m IRT. The values in the OUTA mode was: T.low= -20.0 T.end= 180.0 Tr= 3.0 VOLT= 10. Raised the booms and RTS the instrument at 1939 GMT.

7/16/98

7/17/98

CAM/DC

All tower booms were lowered from 1458 GMT on 7/16/98 until 1513 GMT on 7/17/98 for 25M and 60M T/H calibrations. The booms were powered at ground level during that time.

6/23/98

MW

SDS operations reported a loss of communications with all of the 60M tower instruments beginning at 1004 GMT today. Power cycled the tower COS at 1404 GMT. Communications restored after the power cycle. RTS at 1405 GMT.

5/1/98

CM/MW

The A side 25M and 60M booms were lowered at 1826 GMT for chilled mirror removal. Raised booms and RTS at 1935 GMT.

12/19/97

CAM

Lowered booms from 2100 GMT to 2200 GMT for chilled mirror maintenance.

9/10/97

DN

Lowered the tower from 1915 to 1950 GMT for IOP instrument installation preliminary.

9/04/97

CAM

Scott Richardson was on site and needed to lower the east (A) booms of the tower for planning and preparation of mounting instruments on the booms for the upcoming IOP. Lowered the south (A) 25 and 60 meter booms at 1415 GMT for measurements. Raised booms and RTS at 1445 GMT.

8/14/97

GRS

Raised the tower at 1549 GMT. RTS IRT.

8/12/97

GRS

Lowered the 25 and 60 booms on 8/12/97 at 1625 GMT to perform temp/hum installs and calibrations. The IRT will remain down at ground level until temp/hum work is complete.

4/22/96

CAM/ZW

1810 GMT Lowered tower to T/S IRT. Found Config. For A to D incorrect (mA instead of V). Corrected and tested OK at 1900 GMT.

4/19/96

CAW/CAM

1330 GMT Lowered tower and disconnected IRT at mentors request.

4/18/96

CAW/CAM

1412 GMT Lowered tower to T/S IRT inop. Found loose connection. Corrected and tested OK at 1433 GMT.

4/15/96

CAW/MDW

Installed 25M IRT.